OpenStudy (anonymous):

The volume in cubic feet of a box can be expressed V(x) = x^3 - 6x^2 + 8x, or as the product of three linear factors with integer coefficients. The width of the box is 2 – x. a. Factor the polynomial to find linear expressions for the height and the width. b. Graph the function. Find the x-intercepts. What do they represent? c. Describe a realistic domain for the function. d. Find the maximum volume of the box.

OpenStudy (anonymous):

What's the problem? Factoring, graphing, finding the x-intercepts, finding a domain, or finding maximum volume? ("All of the above" is not an answer, heh)

OpenStudy (anonymous):

\[V(x) = x^3 - 6x^2 + 8x\] \[V(x) = x(x^2-6x+8)\] \[=x(x-2)(x-4)\],

OpenStudy (anonymous):

i guess if one of the factors is supposed to be \(2-x\) you would have to write \[V(x)=-x(x-4)(2-x)\]

OpenStudy (anonymous):

here is a nice graph finding the max is a calculus problem http://www.wolframalpha.com/input/?i=x^3+-+6x^2+%2B+8x++++domain+0..2

OpenStudy (anonymous):

in fact i would say the domain is \((0,1)\) because the sides of the box must be positive

OpenStudy (anonymous):

@satellite73: Please please please don't just DO the problem for people asking questions. It gives no incentive to learn.

OpenStudy (valpey):

Or more likely:? \[V(x)=x(4−x)(2−x)\]

OpenStudy (anonymous):

i guess that would work too right?

OpenStudy (valpey):

Usually in spatial problems we have non-negative dimensions, so it is useful to describe them as such. It makes sense to me that x would be greater than 0 and less than 2.